OpenStudy (anonymous):

find the fundamental matrix e^At for the linear system: x' = [1 1 0] x [0 1 0] [1 0 1]

OpenStudy (anonymous):

i believe that the given matrix is the matrix A in x = e^At , i.e no work required.

OpenStudy (anonymous):

no.. there is some work that needs to be done.

OpenStudy (anonymous):

Did you find the eigenvalues of the given matrix?

OpenStudy (anonymous):

\[\begin{vmatrix} 1-\lambda&1&0\\ 0&1-\lambda&0\\ 1&0&1-\lambda \end{vmatrix} =(1-\lambda)^3=0~\Rightarrow~\lambda_1,\lambda_2,\lambda_3=1\] The eigenvalue \(\lambda=1\) has multiplicity 3, so you'll have to find a generalized eigenvector, if I recall correctly.

OpenStudy (anonymous):

i got to this part.. i don't know how to go further.

OpenStudy (anonymous):

Do you know how to do it using putzer's algorithm? Using e^At?

OpenStudy (anonymous):

I haven't heard of that algorithm, no. But I have learned a thing or two about matrix exponentials.

OpenStudy (anonymous):

okay... can you help me go further please?

OpenStudy (anonymous):

Just a sec... I need to see if I can work this out.

OpenStudy (anonymous):

First eigenvector: Solve for \(\vec{a}\) such that \(A-\lambda I=0\), or equivalently, \[\begin{pmatrix} 0&1&0\\0&0&0\\1&0&0\end{pmatrix}\begin{pmatrix}a_1\\a_2\\a_3\end{pmatrix}=\begin{pmatrix}0\\0\\0\end{pmatrix}\] Solving yields \(a_1=0,a_2=0\) and \(a_3\) can be any real number. Pick any non-zero real for \(a_3\); I'll pick \(1\). So your first eigenvector is \[\vec{a}_1=\begin{pmatrix}0\\0\\1\end{pmatrix}\]

OpenStudy (anonymous):

Have you covered generalized eigenvectors yet? I don't want to present anything new because it's been a while since I've learned all this and I doubt I'd be able to introduce it with ease.

OpenStudy (anonymous):

no we haven't... so using the same method, i find out a2 and a3?

OpenStudy (anonymous):

Not the exact same method, no. It's a bit different. The first generalized eigenvector is obtained by solving for \(\vec{a}_2\) in \((A-\lambda I)\vec{a}_2=\vec{a}_1\). \[\begin{pmatrix} 0&1&0\\0&0&0\\1&0&0\end{pmatrix}\begin{pmatrix}a_1\\a_2\\a_3\end{pmatrix}=\begin{pmatrix}0\\0\\1\end{pmatrix}\] This time, you get \(a_1=1,a_2=0\) and \(a_3\) can be any real. Pick \(a_3=0\) for convenience. You want something that makes \(\vec{a}_1\) and \(\vec{a}_2\) linearly independent. \[\vec{a}_2=\begin{pmatrix}1\\0\\0\end{pmatrix}\]

OpenStudy (anonymous):

The second generalized eigenvector is found similarly. Solve for \(\vec{a}_3\) such that \((A-\lambda I)\vec{a}_3=\vec{a}_2\). \[\begin{pmatrix} 0&1&0\\0&0&0\\1&0&0\end{pmatrix}\begin{pmatrix}a_1\\a_2\\a_3\end{pmatrix}=\begin{pmatrix}1\\0\\0\end{pmatrix}\] This time you get \(a_1=0\) and \(a_2=1\), and \(a_3\) can be any real. Let \(a_3=0\) for ease: \[\vec{a}_3=\begin{pmatrix}0\\1\\0\end{pmatrix}\]

OpenStudy (anonymous):

so is a2 (001)?

OpenStudy (anonymous):

No that's the first eigenvector. \[\vec{a}_1=\begin{pmatrix}0\\0\\1\end{pmatrix},~\vec{a}_2=\begin{pmatrix}1\\0\\0\end{pmatrix},~\vec{a}_3=\begin{pmatrix}0\\1\\0\end{pmatrix}\]

OpenStudy (anonymous):

yeah.. that's what i meant!

OpenStudy (anonymous):

So the linearly independent solutions are \[x_1=e^t\begin{pmatrix}0\\0\\1\end{pmatrix}\] \[x_2=e^t\begin{pmatrix}1\\0\\0\end{pmatrix}\] \[x_3=e^t\begin{pmatrix}0\\1\\0\end{pmatrix}\] giving you the fundamental matrix solution of \[X(t)=\begin{pmatrix}0&e^t&0\\0&0&e^t\\e^t&0&0\end{pmatrix}\] (The eigenvectors multiplied by \(e^t\) make up the columns of the fundamental matrix solution.) Lastly, to find \(e^{At}\), you make use of this theorem: If \(X\) is a fundamental matrix solution of the system \(x'=Ax\), then \(e^{At}=X(t)X^{-1}(0)\).

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

how do you find out a1 again?

OpenStudy (anonymous):

Given some matrix \(A\), you first find the eigenvalues, which are the values of \(\lambda\) such that \(det(A-\lambda I)=0\). Most of the time, you'll you be given a matrix that has non-repeating eigenvalues. In this case, you plug in each of the values of \(\lambda\) and solve the following equation for the corresponding eigenvector: \((A-\lambda I)\vec{v}=\vec{0}\), where \(\vec{v}\) is the corr. eigenvector. Here's a link for more practice: http://tutorial.math.lamar.edu/Classes/DE/LA_Eigen.aspx